# Compression via Pre-trained Transformers: A Study on Byte-Level Multimodal Data

### v1

@@ 6 Discussion
The main goal of our work is to investigate whether pre-trained transformers can be competitive with standard compressors, even when taking their parameter size into account. In contrast to previous work, this places our models into a relatively small regime, where it is unclear whether models will learn well from large datasets at all and have non-trivial out-of-distribution and cross-modality transfer. This could partly be countered by training larger models and then subsequently compressing the model parameters themselves. We chose not to do this in our case since naive lossless compression of model parameters leads to a $10\%$ reduction at best (see Table A3), and even best-case scenarios would only lead to marginal improvements in compression ratio given the size of our largest models. For very large (e.g., foundation) models, compressing weights to achieve competitive compression ratios may be interesting, though it will be necessary to use lossy weight compression techniques (Tao et al., 2022), which lead to non-trivial trade-offs between high (lossy) compression and maintaining strong predictor performance, i.e., the two summands in the numerator of Eq. 2. Exploring these trade-offs is an interesting direction for future research but beyond the scope of our work. Another way to allow for larger models would be to simply evaluate on a larger test set. We deliberately chose to use $1$GB of test data as a regime where standard compression algorithms are hard to beat. Additionally, evaluations on larger test data, and in settings where model parameters are not taken into account have previously conducted (Delétang et al., 2024; Valmeekam et al., 2023; Li et al., 2024) (where significant amounts of cross-domain transfer have also been found, unlike in our experiments).
Note that, similar to Xue et al. (2022), we do not use a tokenizer, which has two reasons. First, tokenizers are typically pre-trained per modality, and we want to rule out bad cross-modality transfer resulting from a bad tokenizer. Second, tokenization acts as a pre-trained “pre-compression” step (Delétang et al. (2024) make a similar comment). This pre-compression increases information density in the context window at the cost of increasing entropy, which can make the prediction problem harder: Lester et al. (2024) even show that when using a strong neural-based pre-compressor (together with arithmetic coding) to train LLMs, training performance can collapse catastrophically.
@@
All our claims regarding the universality of our compressors (or the lack thereof) are limited to the model size regime and the particular modalities and datasets we studied. We cannot rule out that there are cases where even in-modality transfer is weak (e.g., when using another out-of-disribution image evaluation dataset with very different statistics), or that there may be cases of non-trivial cross-modal transfer (which we have not observed). Similarly, our claims regarding outperforming standard compression algorithms are limited to our experiments. We cannot rule out that there are datasets (such as spreadsheet data, or code, which, technically, are both text) where no pre-trained transformer outperforms, e.g., LZMA2 (in fact, we think its plausible that such datasets can be constructed synthetically). Finally, note that the goal of our study is not to build a practical transformer-based universal compressor to compete with standard compressors in terms of computational footprint. As Table A2 shows, our models are orders of magnitude slower for encoding data (and have significantly larger memory- and FLOPS-demands), and they are about three times slower than Bellard’s online adaptive transformer. This is only the forward-pass cost, which can be done for a whole context window at once (without overlap). If our models were used do decode, which has to be performed token-by-token to obtain the correct conditioning, our running time demands would be even worse, making our models clearly uncompetitive in that sense.
@@ 7 Conclusion
In this paper we have shown that it is possible to use pre-trained vanilla transformers as competitive “zero-shot” compressors on out-of-distribution evaluation data, where competitive means achieving better compression ratios than both domain-general and domain-specific standard compression algorithms. We found this to be true for text, images, and audio data, and for all possible combinations of the three — but only as long as the corresponding modalities have been seen during training. We further found that, despite their relatively small size, our models have the capacity to train on multiple modalities, and then compress these well, without losing much performance compared to a purely unimodal model. On the other hand, we found that even multimodal training does not lead to the emergence of a universal compression ability that would yield strong compression performance on unseen modalities. This is in contrast to observations made by Delétang et al. (2024) on LLMs and indicates that there is a qualitative difference between small and (very) large models, even when the small models are trained on large amounts of data. Overall our results suggest that small transformers can be pre-trained to recognize and exploit statistical regularities on par and even better than hand-crafted standard compressors and current state-of-the-art adaptive online neural compressors, but we do not observe the emergence of a general compression ability with our model sizes.
@@ Acknowledgments
We thank Benjamin Beyret, Emilien Dupont, Daniele Calandriello, Grégoire Delétang, Jordi Grau-Moya, Li Kevin Wenliang, Laurent Orseau, Marcus Hutter, Matthew Aitchison, Sarath Chandar, Satinder Baveja, Theophane Weber, Zhengdong Wang, and Zoubin Ghahramani for their helpful feedback and insightful discussions.
Appendix A Experimental Details
A.1 Training Data Sources
@@
A.2 Out-of-Distribution Evaluation Data Sources
@@
We consider a $1$GB subset of the Reddit dataset (Völske et al., 2017), which contains $3.8$ million Reddit posts encoded in UTF-8.
@@ A.3 Sweeps
Fig. 5 in the main paper shows the relationship between context length and model size. For this experiment we performed a large-scale sweep with the goal of covering a good range of training FLOPS budget with models that make various trade-offs between model size and context length (given the same model size, compute demand increases with increasing context length). The main question was whether there is a qualitatively similar relationship across parameters, and whether there is a clear sweet spot — see the main paper for results and discussion. For our sweep we used the same model parameters as in the previous paragraph (the training data size was always at $100\%$) and sweep over the following four context sizes (with training batch size in brackets): ${[1024\leavevmode\nobreak\ (128),2048\leavevmode\nobreak\ (64),4096\leavevmode\nobreak\ (32),8192\leavevmode\nobreak\ (16)]}$. For each context size we train five models (XS, S, M, L, and XL) on all three unimodal datasets, respectively. Each model has a different combination of embedding dimension and number of layers for each different context size. The XS models have embedding dimensions ${[112,96,80,64]}$ and numbers of layers ${[11,7,5,3]}$ for the different context sizes respectively (i.e., wider and deeper models for shorter contexts and more narrow and more shallow models for long context size). The S models have embedding dimensions ${[192,160,112,96]}$ and numbers of layers ${[10,8,6,4]}$. The M models have embedding dimensions ${[224,192,144,112]}$ and numbers of layers ${[12,9,7,5]}$. The L models have embedding dimensions ${[272,240,176,144]}$ and numbers of layers ${[13,10,8,5]}$. The XL models have embedding dimensions ${[320,304,240,160]}$ and numbers of layers ${[12,9,7,6]}$. The main goal with these settings is to create families of models that have roughly the same demand in terms of FLOPS (iso-FLOPS) but very different trade-offs in terms of model- and context size.
A.4 Computational Resources
@@
Appendix B Additional Results
B.1 Compression Ratios
@@
Evaluation Modality Ours Bellard gzip LZMA2 FLAC PNG JPEG 2000
Audio 0.487 0.509 0.813 0.699 0.538 - -
Image 0.285 0.281 0.698 0.545 - 0.426 0.390
Text 0.217 0.204 0.394 0.286 - - -
Audio + Image 0.393 0.395 0.756 0.622 - - -
Audio + Text 0.362 0.357 0.604 0.493 - - -
Image + Text 0.270 0.243 0.546 0.415 - - -
Audio + Image + Text 0.349 0.331 0.635 0.510 - - -
@@
B.2 Running Times
Table A2: Running times to compress $1$GB of data for all compressors used in our study. Note that we use the best model per modality, which have different sizes and thus different running times.
@@ B.2 Running Times
Evaluation Modality Ours Bellard gzip LZMA2 FLAC PNG JPEG 2000
Audio 305 609 101 178 55 524 169 - -
Image 222 065 103 391 47 436 174 495 99
Text 452 355 100 657 102 881 184 - -
Table A2 shows the wall-clock running times in seconds for compressing $1$GB of data from each of the three modalities for our models, Bellard’s online adaptive transformer (Bellard, 2021), and the standard compression algorithms used in our work. As the table clearly shows, our models and Bellard’s model are orders of magnitudes slower (let alone the increased computational demand and GPU requirements). Note that running times for our models differ, because we pick the best model per modality, which are models of different sizes.
B.3 Compressing Model Parameters
@@
Throughout our paper we report compression rates that take uncompressed model parameters into account. As discussed in the main paper, compression ratios could be improved by also compressing model parameters. However, as Table A3 shows, naively compressing model parameters with a lossless compressor does not lead to much compression, which would translate into very marginal gains on the overall compression ratio. While it is possible to investigate more sophisticated compression schemes, in particular lossy compression of network weights (though this opens the problem of having to solve a trade-off between increasing weight compression and maintaining compression performance), this is beyond the scope of our paper. Accordingly, our compression rates can be understood as (somewhat) conservative estimates that give (in our case fairly tight) upper bounds on compression performance. The topic of compressing network weights to achieve competitive compression ratios would be of greater significance in a regime where models are significantly larger than ours (but the evaluation data stays roughly at the same size).
B.4 Scaling Analysis for Multimodal Training
Figure A1: Similar to Fig. 4 in the main paper, but here the models are trained on a uniform mixture over all three modalities ($55$GB per modality). The plot shows compression performance evaluated on the unimodal datasets as training progresses for various model- and training set sizes (models are different colors, each line is a different training set size of either $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$). We always train for $2$ epochs, regardless of dataset size, i.e., smaller datasets require fewer FLOPS. In contrast to Fig. 4, where models are trained on unimodal data, we observe no overfitting, e.g., on images, even for the largest models tested. Note, however, that the compression ratios are slightly worse than for unimodal training, which is in line with our other expriments that show small losses when training on multimodal data.

### v2

@@ 6 Discussion
Our main goal is to investigate whether pre-trained transformers can compete with standard compressors, even when taking their parameter size into account. In contrast to previous work, this places our models into a relatively small regime, where it is unclear whether models will learn well from large datasets at all and have non-trivial out-of-distribution and cross-modality transfer. One could try to train larger models and compress the model parameters themselves, but we chose not to do this since naive lossless compression of model parameters leads to $\leq 10\%$ reduction (see Table A4), and even best-case scenarios would only lead to marginal improvements in compression ratio given the size of our models. For very large (e.g., foundation) models, compressing weights to achieve competitive compression ratios may be interesting, though it would require lossy compression (Tao et al., 2022), leading to non-trivial trade-offs between high (lossy) compression and maintaining strong predictor performance (the two summands in the numerator of Eq. 3). Exploring these trade-offs is an interesting direction for future research but beyond the scope of our work. Another option for larger models is to evaluate on more test data. We chose $1$GB of test data as a regime where standard compression algorithms are hard to beat. Moreover, evaluations on more test data, and/or without accounting for model parameters, have been conducted (Delétang et al., 2024; Valmeekam et al., 2023; Li et al., 2024) (finding significant cross-domain transfer, unlike our experiments).
Like Xue et al. (2022), we do not use a tokenizer, which has two reasons. First, tokenizers are typically pre-trained per modality, and we want to rule out bad cross-modality transfer resulting from a bad tokenizer. Second, tokenization acts as a “pre-compression” step (cf. Delétang et al. (2024)). This pre-compression increases information density in the context window at the cost of increasing entropy, which can make the prediction problem harder: Lester et al. (2024) even show that when using a strong neural-based pre-compressor (together with arithmetic coding) to train LLMs, training performance can collapse catastrophically.
@@
All our claims regarding the universality of our compressors (or the lack thereof) are limited to the model size, regime, and the particular modalities and datasets we studied. We cannot rule out that there are cases where even in-modality transfer is weak (e.g., using another OOD image evaluation dataset with very different statistics), or that there are cases of non-trivial cross-modal transfer (which we have not observed). We did not investigate transfer learning approaches to improve the out-of-modality performance of our neural compressors, but we consider this an interesting avenue for future work. Our claims regarding outperforming standard compression algorithms are limited to our experiments. We cannot rule out that there are datasets where no pre-trained transformer outperforms, e.g., LZMA2 (in fact, we think its plausible that such datasets can be constructed synthetically). Moreover, we cannot rule out that more sophisticated architectures (e.g., Perceivers (Jaegle et al., 2021), MegaByte (Yu et al., 2023), or Byte Latent Transformers (Pagnoni et al., 2024)), would outperform our models, and we consider neural architecture optimization for lossless compression an interesting direction for future research. Finally, our goal is not to build practical transformer-based universal compressors to compete with standard compressors in terms of computational footprint. As Table A3 shows, our models are orders of magnitude slower for encoding data (with significantly larger memory- and FLOPS-demands) and $\approx 3$x slower than Bellard’s online adaptive transformer. When decoding data with our models, which has to be done token-by-token to obtain the correct conditioning, our running time is even worse.
Future Work
Even our largest models that were trained on multiple different modalities failed to achieve out-of-modality transfer (unlike LLMs, which do achieve cross-modal transfer when used as compressors, albeit using the “unajdusted” compression ratio; see Delétang et al. (2024)). Accordingly, investigating more sophisticated training strategies to improve cross-modal transfer, e.g., transfer learning, presents an interesting direction for future work.
@@ 7 Conclusion
We showed that pre-trained vanilla transformers are competitive “zero-shot” compressors on OOD evaluation data, achieving better compression ratios than domain-general and domain-specific standard compression algorithms. We showed this for text, images, and audio data, and for all possible combinations of the three — as long as the corresponding modalities are in the training mixture. Despite their relatively small size, our models compress well across multiple modalities, without losing much performance compared to a purely unimodal model. On the other hand, multimodal training does not lead to strong compression performance on unseen modalities. This is in contrast to LLMs (Delétang et al., 2024), indicating a qualitative difference between small and (very) large models, even when the small models are trained on large datasets. Our results suggest that small transformers can be pre-trained to exploit statistical regularities on par or better than standard compressors and current state-of-the-art adaptive online neural compressors, but we do not observe the emergence of universal compression.
Impact Statement
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgments
We thank Anna Mitenkova, Benjamin Beyret, Emilien Dupont, Daniele Calandriello, Grégoire Delétang, Jordi Grau-Moya, Li Kevin Wenliang, Laurent Orseau, Marcus Hutter, Matthew Aitchison, Sarath Chandar, Satinder Baveja, Theophane Weber, Zhengdong Wang, Zoubin Ghahramani, and the anonymous reviewers for their helpful feedback and insightful discussions.
Appendix A Relation to Prior Work
Table A1: Similarities and differences of our work w.r.t. Bellard (2021) and Delétang et al. (2024). Unlike Bellard (2021), we use offline (i.e., pre-) training and consider multimodal data. Online training imposes different constraints on the optimal model size, since the compression ratio does not have to account for the model parameters, which is why Bellard (2021) were able to study somewhat larger models. We also conduct comprehensive scientific ablations over dataset-, model-, context-size, etc. Unlike Delétang et al. (2024), we (only) consider the adjusted (i.e., the actual) compression ratio, which includes the model parameters, and therefore, we cannot use massive foundation models. Moreover, while Delétang et al. (2024) studied the cross-domain transfer of foundation models trained on text, they did not study multimodal training.
Bellard (2021) Delétang et al. (2024) Ours
# of Parameters $187$M $70$B $20$M
Training online offline offline
Multimodal - evaluation training & evaluation
Actual (Adj.) Compression Ratio
Ablations Model Size
Context Size
Tokenizer
Data Mixture
Train Dataset Size
Evaluation Dataset Size
Sliding Window Overlap
As stated in Section 1 and shown in Table A1, we study the open question of whether small transformers pre-trained on multimodal data can achieve competitive compression ratios across different modalities and whether there is transfer to unseen modalities (as observed in the large-scale model case). Consequently, the fact that they can do so cannot be predicted from the strong compression performance of LLMs (e.g., results by Delétang et al. (2024)).
In contrast to Bellard (2021), we use offline adaptive arithmetic coding, which induces entirely different constraints on the optimal model size (since the model parameters do not have to be factored into the compression ratio). Bellard (2021) relies on “test-time” gradient updates on in-distribution data, whereas we pre-train models and then leverage in-context learning on out-of-distribution data. It is, therefore, impossible to conclude from the results in Bellard (2021) whether our results are possible. The training data and protocol are incomparable, making it all the more interesting that our results are generally very close — another surprising finding that cannot be trivially explained.
In contrast to Delétang et al. (2024), we do not evaluate off-the-shelf, large-scale, text-based foundation models but pre-train small-scale transformers on audio, image, and text data. As a result, the compressors proposed by Delétang et al. (2024) are not competitive w.r.t. standard compressors (which they acknowledge) — unlike our models (we beat standard compressors across the board). Delétang et al. (2024) do conduct a pilot experiment by pre-training a small transformer on text data, but, in contrast to our work, they do not perform (i) a comprehensive study of multi-modal training, (ii) out-of-distribution evaluation, (iii) ablations over context-, dataset-, and model-size, (iv) a comparison to Bellard (2021), (v) an investigation of the sliding window overlap, and (vi) a test dataset size ablation. As our work shows, all of these ingredients are necessary to obtain the strong performance we reported, and, as a result, Delétang et al. (2024) do not manage to pre-train a transformer-based compressor that beats standard compressors across multiple modalities — unlike our work. The main contribution of Delétang et al. (2024) is conceptual, with a relatively simple experimental evaluation to illustrate the conceptual arguments. Our work, on the other hand, performs a rigorous and comprehensive empirical study.
Appendix B Experimental Details
B.1 Training Data Sources
@@
B.2 Out-of-Distribution Evaluation Data Sources
@@
We use a $1$GB subset of Reddit (Völske et al., 2017), which contains $3.8$ million Reddit posts encoded in UTF-8.
@@ B.3 Sweeps
Fig. 5 in the main paper shows the relationship between context length and model size. For this experiment we performed a large-scale sweep with the goal of covering a good range of training FLOPS budget with models that make various trade-offs between model size and context length (given the same model size, compute demand increases with increasing context length). The main question was whether there is a qualitatively similar relationship across parameters, and whether there is a clear sweet spot — see the main paper for results and discussion. For our sweep we used the same model parameters as in the previous paragraph (the training data size was always at $100\%$) and sweep over the following four context sizes (with training batch size in brackets): ${[1024~{}(128),2048~{}(64),4096~{}(32),8192~{}(16)]}$. For each context size we train five models (XS, S, M, L, and XL) on all three unimodal datasets, respectively. Each model has a different combination of embedding dimension and number of layers for each different context size. The XS models have embedding dimensions ${[112,96,80,64]}$ and numbers of layers ${[11,7,5,3]}$ for the different context sizes respectively (i.e., wider and deeper models for shorter contexts and more narrow and more shallow models for long context size). The S models have embedding dimensions ${[192,160,112,96]}$ and numbers of layers ${[10,8,6,4]}$. The M models have embedding dimensions ${[224,192,144,112]}$ and numbers of layers ${[12,9,7,5]}$. The L models have embedding dimensions ${[272,240,176,144]}$ and numbers of layers ${[13,10,8,5]}$. The XL models have embedding dimensions ${[320,304,240,160]}$ and numbers of layers ${[12,9,7,6]}$. The main goal with these settings is to create families of models that have roughly the same demand in terms of FLOPS (iso-FLOPS) but very different trade-offs in terms of model- and context size.
B.4 Computational Resources
@@
Appendix C Additional Results
C.1 Compression Ratios
@@
Evaluation Modality Ours Bellard gzip LZMA2 FLAC PNG JPEG 2000 JPEG-XL
Audio 0.487 0.509 0.813 0.699 0.538 - - -
Image 0.285 0.281 0.698 0.545 - 0.426 0.390 0.260
Text 0.217 0.204 0.394 0.286 - - - -
Audio + Image 0.393 0.395 0.756 0.622 - - - -
Audio + Text 0.362 0.357 0.604 0.493 - - - -
Image + Text 0.270 0.243 0.546 0.415 - - - -
Audio + Image + Text 0.349 0.331 0.635 0.510 - - - -
@@
C.2 Running Times
Table A3: Running times to compress $1$GB of data for all compressors used in our study. We use the best model per modality, which have different sizes and thus different running times.
@@ C.2 Running Times
Evaluation Modality Ours Bellard gzip LZMA2 FLAC PNG JPEG 2000 JPEG-XL
Audio 305 609 101 178 55 524 169 - - -
Image 222 065 103 391 47 436 174 495 99 871
Text 452 355 100 657 102 881 184 - - -
Computing the FLOPS for standard compressors (e.g., gzip or PNG) is much more involved than for neural networks and requires intricate knowledge of the algorithm, which can consist of more than 100K lines of code (standard compression algorithms have also been highly optimized and run on CPU, whereas the neural compressors need a GPU). Therefore, we instead compare the wall-clock running times in seconds when compressing $1$GB of data from each of the three modalities for our models, Bellard’s online adaptive transformer (Bellard, 2021), and the standard compression algorithms used in our work. As Table A3 clearly shows, our models and Bellard’s model are orders of magnitudes slower (let alone the increased computational demand and GPU requirements). Running times for our models differ, because we pick the best model per modality, which are models of different sizes.
C.3 Compressing Model Parameters
@@
Throughout our paper we report compression rates that take uncompressed model parameters into account. As discussed in the main paper, compression ratios could be improved by also compressing model parameters. However, as Table A4 shows, naively compressing model parameters with a lossless compressor does not lead to much compression, which would translate into very marginal gains on the overall compression ratio. While it is possible to investigate more sophisticated compression schemes, in particular lossy compression of network weights (though this opens the problem of having to solve a trade-off between increasing weight compression and maintaining compression performance), this is beyond the scope of our paper. Accordingly, our compression rates can be understood as (somewhat) conservative estimates that give (in our case fairly tight) upper bounds on compression performance. Compressing network weights to achieve competitive compression ratios would be of greater significance in a regime where models are significantly larger than ours (but the evaluation data stays roughly at the same size).
C.4 Scaling Analysis for Multimodal Training
Figure A1: Similar to Fig. 4 in the main paper, but here the models are trained on a uniform mixture over all three modalities ($55$GB per modality). The plot shows compression performance evaluated on the unimodal datasets as training progresses for various model- and training set sizes (models are different colors, each line is a different training set size of either $20\%$, $40\%$, $60\%$, $80\%$, and $100\%$). We always train for $2$ epochs, regardless of dataset size, i.e., smaller datasets require fewer FLOPS. In contrast to Fig. 4, where models are trained on unimodal data, we observe no overfitting, e.g., on images, even for the largest models tested. However, the compression ratios are slightly worse than for unimodal training, which is in line with our other expriments that show small losses when training on multimodal data.